How would I do these? Please show all work
Kason a sports yiths 1e 2012 NBA champldnchip?
Answer #1

#1.
first number can be selected in 10C1 ways
second can be selected in 9C1 ways as repetation is not allowed
third in 8C1 and fourth in 7C1 ways

Hence total possible opening combinations = 10*9*8*7 = 5040

#2.
Odds of the lakers winning = (1/101)/(1 - 1/101) = (1/101)/(100/101)
= 1/100